OpenStudy (anonymous):

write the expression in factored form? x^2+8x-xy-8y

OpenStudy (anonymous):

Somebody please help me:)!!!!!!

OpenStudy (anonymous):

try grouping like this first: \(\large x^2+8x-xy-8y \) = \(\large (x^2+8x)-(xy+8y) \) now do you see what you can factor our from each of the groupings (parenthesis) ????

OpenStudy (anonymous):

*factor out

OpenStudy (anonymous):

am i suppose to still work on that or is that the answer

OpenStudy (anonymous):

i gave a hint so you can continue...

OpenStudy (anonymous):

what else am i suppose to do to get the answer please help

OpenStudy (anonymous):

look at the first grouping, \(\large (x^2+8x) \) do you see what you can factor out here? each term in this binomial has the common factor of x

OpenStudy (anonymous):

will it be (x^2-1) ?

OpenStudy (anonymous):

i dont know im so stupid ughh

OpenStudy (anonymous):

no... notice that \(\large (x^2+8x)=(x\cdot \color{red}x+8 \cdot \color{red}x)=(x+8)\cdot \color{red}x \) ok... i factored the fist term for you... can you do the same for the second grouping: \(\large (xy+8y) \)

OpenStudy (anonymous):

i tried working on it a bit now and i got something like x(x+8)

OpenStudy (anonymous):

ok.... now do the second grouping from above....

OpenStudy (anonymous):

so far i have -x(y-8) but that doesnt look right

OpenStudy (anonymous):

no, it's not... what do both xy and 8y have in common?

OpenStudy (anonymous):

y? so itll be y(x+8)?

OpenStudy (anonymous):

correct... :)

OpenStudy (anonymous):

ok i have that now how do i write my answer out?

OpenStudy (anonymous):

one more step. so this is what we have so far: \(\large x(x+8)-y(x+8) \) look at the big picture... notice they both have (x+8) as a common factor... factor this whole thing out just as you did before...

OpenStudy (anonymous):

i dont know how to factor it out.? can you help me:x

OpenStudy (anonymous):

suppose I give you this as an example: \(\large xa-ya \) how would you factor this?

OpenStudy (anonymous):

x(a) - y(a) ??? i really dont know how to do this and i even did the tutorial alot of times and still dont get it

OpenStudy (anonymous):

both terms have "a" in common so factor that out: (x - y)a do you understand this part?

OpenStudy (anonymous):

could it be (x-y)( x+8)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yayyy thankss now i have like 9 more to go ughhh